Bwindi Impenetrable Forest (04.01.2011 to 06.30.2011)

You are here

PDF versionPDF version
April - June
Communications activities: 

No workshops at site

Unusual Events at the TEAM Site: 

No unusual events at the site

New Species: 

No new species found or identified at the site

Protocol Activities: 

30 camera traps for array one where set and retrieved in this quarter. Setting for aray two is in progress.

Protocol Problems: 

Power shortages at the climate station were experianced. The problem was solved by adding a second solar panel. Power problems are still being experienced. A second battery yet to be installed. The problem was communicated and discussed with the technical director.

Schedule Problems: 

No deviations on the schedule. All protocols are on schedule

Logistical Problems: 

No logistical problems.


No suggestions for TEAM technical/ administrative staff